YEAR 1 (R-Billion) YEAR 2 (R-Billion)

Investment 200 220



Saving 180 190

Export 100 110

Imports 120 140

Government Expenditure 150 160

Taxation 150 160

Equilibrium National Income 1 800 2 000



(a) Calculate the value of the multiplier for this economy.

(b) Should the full employment level of income be R-B2 255, by how much should government change its spending to reach this level of income during the next year given the multiplier has not changed.

(c) Evaluate whether or not this policy approach is effective in real life in achieving the desired level of GDP.
1
Expert's answer
2021-07-27T04:46:01-0400

(a) value of the multiplier

relationship between total national income and government spending.

"=\frac{1800}{150}=12" for year 1


"=\frac{2000}{160}=12.5" for year 2

This implies that the value of the multiplier is "12\times" .

(b)

level of income=2255

multiplier= "12\times"

"12=\frac{2255}{x}"

"12x=2255"

"x=\frac{2255}{12}"

"x=187.9"

this is approximately 188.

the government should thus increase its spending by a value of (188-160) which is 28.

(c)

This policy is effective in achieving the desired level of GDP . This is because it relates the equilibrium national level of income directly to the government expenditure.